Given: Neither a nor b equal zero.
Evaluate:
(a+bi)[(a/a^2+b^2)-(b/a^2+b^2)i]

Answer by stanbon(75887)

Given: Neither a nor b equal zero.
Evaluate:
(a+bi)[(a/a^2+b^2)-(b/a^2+b^2)i]
---------------------------------------
Since a^2+b^2 = (a+bi)(a-bi) you get the following:
---
= a/(a-bi) - b/(a-bi)
---
= (a-b)/(a-bi)
====================
Cheers,
Stan H.